OpenStudy (moonlitfate):

Find the derivative of (x^4)*(3-9x)^(1/2)

OpenStudy (moonlitfate):

I know that the power rule has to be used, and I think the chain rule, but I'm still not the greatest at the latter. :/

OpenStudy (moonlitfate):

I know that the derivative of x^4 = 4x^3...

OpenStudy (moonlitfate):

The derivative of (3-9x)^(1/2) = 1/2 (3-9x)^(-1.2), I think...

OpenStudy (anonymous):

you are going to take lots of derivatives so first off just memorize this: \[\frac{d}{dx}\sqrt{x}=\frac{1}{2\sqrt{x}}\] it never changes, so forget about using the power rule with \(\frac{1}{2}\) just straight up remember it. that way when your classmates are all converting to exponents, using the power rule, and converting back, you are already done. like memorizing seven times eight

OpenStudy (moonlitfate):

I still don't have 7*8 memorized. xD But, I'll make sure to memorize that, it just helps.

OpenStudy (anonymous):

so by the chain rule, the derivative of \[\sqrt{3-9x}\] is \[\frac{1}{2\sqrt{3-9x}}\times (-9)=\frac{-9}{2\sqrt{3-9x}}\]

OpenStudy (anonymous):

yeah i know, that is why i use seven times eight as an example all the others everyone knows, that you just have to memorize

OpenStudy (anonymous):

of course for \[x^4\sqrt{3-9x}\] you need both the product rule and the chain rule

OpenStudy (moonlitfate):

So, it would be the derivative of\[x^4\sqrt{3-9x} * \frac{ -9 }{ 2\sqrt{3x-9} }\] ; not the ddrivative of the whole thing, I just the first function * what was all ready found out

OpenStudy (moonlitfate):

Chain rule is mind boggling business.

OpenStudy (anonymous):

you will also need to use the product rule.

OpenStudy (anonymous):

f = x^4 f' = 4x^3 g= (3-9x)^1/2 g' = -9/2(sqrt (3-9x) f'*g+g'*f

OpenStudy (moonlitfate):

Thanks, I'll see what I can do from there. :)